# Dynamics and fluid-structure interaction in turbulent flows within and above flexible canopies

### 21 Dec 2023 v1

@@ 3.5 Filament density variation
In order to investigate the effects of a variation in the inertial properties of the filaments, we consider two additional values of the density ratio $\{1.0+1.46\cdot 10^{-2},1.0+1.46\cdot 10^{-1}\}$ starting from the case at $Ca=25$ and $\rho_{s}/\rho_{f}=1.0+1.46\cdot 10^{-3}$; $Ca$ is left unchanged, as well as all the other simulation parameters. We evolve the simulations until the attainment of a statistically steady state and investigate it with the same tools introduced earlier.
@@ 3.6 The effect of the filaments’ flexibility
The drag reduction observed freezing a flexible canopy in its instantaneous configuration can therefore be imputed to an overall reduction of the turbulent activity with respect to the flexible case. We identify two phenomena contributing to such effect: first, upon freezing the filaments, no more energy is pumped into the small scale turbulent fluctuations of the flow by their flapping motion. This is consistent with the maximum drag reduction being attained in the intermediate regime ($Ca\approx 25/50$), where the large amplitude flapping of the filaments is most regular. As further discussed in the following section, in facts, at intermediate values of $Ca$ the natural response of the filaments resonates with a characteristic frequency of the turbulent forcing. Second, the frozen canopies are more effective then the flexible ones at preventing intense vertical interactions between the inner and the outer flows, thus yielding a weakly driven turbulent state within the canopy and depleting the turbulent fluctuations close to the shear layer.
@@ 4.1 Characterisation of motion
Multiple quantities can be observed in order to characterise the motion of the flexible filaments constituting the canopy. Here, we focus on the spanwise velocity of the Lagrangian points at the tip of the filaments: we prefer the measurement of a velocity to that of a displacement since it is more directly related to the turbulent fluctuations of the flow, and we favour the spanwise direction as it is the one less affected by the constraints of the problem. The streamwise and vertical dynamics of the filaments are the most influenced by the constraints, i.e., the inextensibility and the presence of the wall. We therefore collect the Lagrangian velocity signal in the spanwise direction at the tip of all the filaments throughout all the values of $Ca$ considered in this study, and highlight its features both in the time and frequency domains. For the purpose of visualisation, we also choose one canopy filament per value of $Ca$ and show in panel $a$ of figure 17 the velocity signal over ten bulk time units extracted after the attainment of a fully developed flow state. Here we introduce the results of an additional simulation at $Ca=500$ to encompass the case in which the filaments are fully compliant to the flow.
@@
We immediately observe that the amplitude of the fluctuations increases with the flexibility of the filaments up to $Ca=25/50$, where it appears to saturate. Furthermore, the small amplitude fluctuations of the most rigid case ($Ca=1$) are characterised by a nearly sinusoidal shape, modulated in amplitude by slower dynamics. Such behaviour contrasts with the motion of the most flexible filaments ($Ca=100/500$), which instead exhibit fluctuations with a less definite shape and period. In these cases it is also possible to appreciate abrupt and fast small scale fluctuations on top of the slower dynamics: those arise from the collisions of the filament with the wall, and indeed are not observed for the lowest values of $Ca$ as the filament never gets sufficiently deflected. In between these two behaviours ($Ca=50$), the large amplitude fluctuations are almost sinusoidal in shape and exhibit a more definite period; they therefore appear compatible with a resonance between the natural structural response of the filaments (responsible for the nearly sinusoidal oscillations of the most rigid cases) and the turbulent fluctuations of the flow (dominating the motion of the filaments in the most flexible cases).
@@
To confirm the picture above, we investigate the energy spectrum of the signal, $E_{z}^{tip}=\frac{1}{2}\overline{\overline{\mathcal{F}(w_{tip})\mathcal{F}^{*}(w_{tip})}}$, where $\mathcal{F}$ is the temporal Fourier’s transform operator, *** is the conjugate of an arbitrary complex number and the double over bar denotes ensemble-averaging on all the filaments of each case. After computing $E_{z}^{tip}$ over a sufficiently long time span cleaned of any initial transient, we report it in panel $b$ of figure 17 for the different values of $Ca$ considered. We also mark with vertical dashed lines the natural frequencies associated to the first structural mode of the different filaments, $f_{nat}$, computed as described in §2. For the lowest values of $Ca$ ($1$ to $50$), the spectrum peaks in correspondence of the natural frequency (and higher harmonics), thus confirming that the structural natural response dominates the dynamics of the most rigid filaments. The value of $f_{nat}$ decreases with $\gamma$, hence it gets shifted towards the left for increasing values of $Ca$. The peak of the spectrum also obeys this shift in frequency up to $Ca=50$, while beyond that its position does not to change significantly with the $Ca$. Such behaviour suggests that the most flexible filaments adjust to a motion no more dictated by their structural properties, but presumably compliant to the turbulent fluctuations of the flow. The transition between the two distinct regimes of motion occurs at $Ca\approx 50$, and thus we justify the large amplitude nearly sinusoidal oscillations observed in that case with the resonance between the first structural mode of the filaments and a characteristic frequency of the turbulent forcing to which they are exposed.
@@
The identification of two distinct regimes of motion is not unexpected: previous investigations addressing the motion of flexible fibres in homogeneous isotropic turbulence (Rosti et al., 2018a, 2020; Olivieri et al., 2022b) highlighted the existence of a structure-dominated and a turbulence-dominated regime, partially similar to what found when studying vortex induced vibrations (Bearman, 1984). We recently characterised numerically the motion of a flexible fibre clamped in wall turbulence over a wide range of structural parameters (Foggi Rota et al., 2023), also observing the emergence of the two regimes discussed above. Recently, Fu et al. (2023) experimentally investigated the individual motion of an isolated flexible plant model and how that changes when multiple plants are arranged in a canopy, while Monti et al. (2023) studied the motion of the filaments in a flexible canopy, mostly to characterise their collective dynamics (honami/monami). To reconcile the different insights offered by all these works, in the following section we further describe the flapping state of the individual filaments and better contextualise our observations.
@@ 4.2 The flapping state
To better investigate the different dynamical regimes of the filaments, we compute their dominant flapping frequency along the spanwise direction, $f^{flap}_{z}$, i.e., the peak location of the spectra reported in panel $b$ of figure 17, and plot it against the natural frequency associated to the corresponding value of $Ca$, i.e., $f_{nat}\approx 0.45\frac{U_{b}}{H}\sqrt{\frac{\rho_{f}}{\rho_{s}}\frac{h}{d}\frac{1}{Ca}}$. Since for the highest values of $Ca$ the position of the spectral peak does not appear to undergo any significant change with the structural parameters of the filaments, we denote it as $f_{turb}$ (i.e., the characteristic flapping frequency of the filaments in the turbulence dominated regime) and employ it to adimensionalise both $f^{flap}_{z}$ and $f_{turb}$. The outcome is reported in panel $a$ of figure 18, where flapping frequencies close to the natural one lay on the dashed-dotted line with unitary slope, $f^{flap}_{z}/f_{turb}=f_{nat}/f_{turb}$, while those close to the chosen value of $f_{turb}$ approach the dashed horizontal asymptote. We thus confirm that the most rigid filaments exhibit their natural response, while the most flexible ones deviate from it and approach $f^{flap}_{z}\approx f_{turb}=0.5U_{b}/H$. Such frequency, independent from the structural characteristics of the filaments, is consistent with the outcome of our previous investigation (Foggi Rota et al., 2023) focused on the flapping states of an isolated flexible fibre clamped in wall turbulence. Also in that case indeed, we identified a turbulence-dominated regime for the highest values of $Ca$ where the fibre is compliant to the flow and its motion is characterised by a slow flapping with a period comparable to the bulk time-scale of the flow. Furthermore, despite the absence of a physical justification for $f^{flap}_{z}\approx 0.5U_{b}/H$, we observe that its occurrence in both the motion of an isolated flexible fibre in a full-channel flow and in the motion of the canopy filaments in the the half-channel considered here supports the reproducibility and generality of the result.
@@
Figure 18: Flapping states of the filaments for different values of $Ca$, hence $f_{nat}/f_{turb}$. We first show (panel $a$) the dominant frequency of oscillation along the spanwise direction, $f^{flap}_{z}$, extracted as the peak location of the spectra in figure 17. Squares denote data from the cases at different density ratios. (panel $b$) We also report the trends of the elastic ($\Gamma_{s}$) and kinetic ($K_{s}$) energy of the filaments per unit length, averaged over time and over all the filaments.
@@
In order to assess the effect of an increase in the structural density of the filaments on their dynamics, we also report in the map of figure 18, panel $a$, the results of the cases at ${\rho_{s}}/{\rho_{f}}=\{1+1.46\cdot 10^{-2},1+1.46\cdot 10^{-1}\}$, respectively associated to a higher and a lower value of the natural frequency. Differently from what found in our previous investigation of a single filament (Foggi Rota et al., 2023), here an increase in ${\rho_{s}}/{\rho_{f}}$ prevents the filaments from flapping at $f_{turb}$ and they thus exhibit their natural response. This difference originates from the fact that denser filaments yield a stronger obstruction of the mean flow in the case of a canopy (as elucidated in §3.5), and consequently get less deflected and maintain an almost upright configuration, swaying at $f_{nat}$. This effect is lost in the case of an isolated fibre, as the back-reaction on the mean flow is negligible; in this isolated case the fibre is less shielded and consequently gets more deflected by the flow, swaying with it.
@@
The effect of the clamp on the dynamics of the filaments is further understood observing the variation of their elastic energy with $Ca$, while their motion is better described in terms of kinetic energy. To those ends we therefore consider the weak form of equations 3 and 4, integrated in time and over the length of the filaments
@@
represents the structural kinetic energy per unit length, likewise averaged. We report the trends of $\Gamma_{s}$ and $K_{s}$ with $f_{nat}/f_{turb}$, and hence with the $Ca$, in panel $b$ of figure 18, where the error-bars lay within the size of the markers. $\Gamma_{s}$ peaks well within the regime dominated by the natural response of the filaments, contrasting with the behaviour of a free fibre addressed by Rosti et al. (2018a), but confirming what can also be observed in the case of an isolated clamped fibre. We therefore deduce that the clamp to the wall, constraining the shape of the filaments, hinders their deflection and shifts the maximum of $\Gamma_{s}$ towards higher values of $\gamma$. The kinetic energy, instead, follows a trend consistent with the amplitude of the fluctuations in figure 17, panel $a$, increasing with the $Ca$ in the most rigid cases and reaching a plateau close to the resonance between $f_{nat}$ and $f_{turb}$.
Characterising the dominant dynamics of our canopy filaments, we have been able to compare our results with those attained considering a single fibre clamped in wall turbulence (Foggi Rota et al., 2023) and plants with a different geometry (Fu et al., 2023). We have also isolated the role played by the sheltering effect of the canopy upon increasing the density of the filaments, and by the model adopted to describe them. Nevertheless, we emphasise that the present investigation is unable to capture any sub-dominant dynamics of the filaments, such as the contribution of turbulence to their motion in the regime dominated by the natural response. Regardless of the value of $Ca$, in facts, the coherent motion of the filaments (honami/monami) investigated by Monti et al. (2023) remains dictated by the turbulent structures in the flow.
@@ 5 Conclusions
Filaments of increasing flexibility are more deflected by the flow, thus reducing the frontal area of the canopy, and they yield a weaker drag discontinuity at the canopy tip, associated to velocity fluctuations less intense then in the rigid case. Consequently, the canopy drag decreases increasing the flexibility up to saturation, attained when the fully deflected filaments start piling on top of each other parallel to the wall. The spectra of the turbulent kinetic energy within the canopy confirm the spectral short cut mechanism described by Finnigan (2000) and Olivieri et al. (2020), while a more regular energy cascade is attained above the canopy tip. Furthermore, the intensity of the velocity fluctuations within the canopy increases with its flexibility due to the increased motion of the filaments. Outside the canopy, instead, the opposite trend is attained by virtue of the generation of a stronger shear layer in the more rigid cases.
The characterisation of the turbulent state at different distances from the bottom wall suggests a multi-layer approach to turbulence modelling in canopy flows. Within the canopy, turbulence is quasi 2D close to the bottom wall and it approaches an isotropic state between the inner inflection point and the virtual origin of the outer flow. The picture becomes more blurred increasing the flexibility, as the flow in the canopy turns less isotropic. A strongly anisotropic state is attained approaching the canopy tip, and the flow exhibits a behaviour consistent with what observed in conventional channel flows moving above it, thus supporting outer similarity arguments. The turbulent flow above the canopy appears populated by spanwise vortical structures likely generated by the drag discontinuity at the canopy tip, consistently with what observed in previous investigations (Finnigan, 2000). Those structures induce sweep and ejection events dominating the interaction between the inner and the outer flows, but only the strongest sweep events are able to penetrate the canopy due to the obstruction exerted by the deflected filaments. The effect increases with the flexibility, yielding a more weakly driven turbulent state within the most compliant canopies even though the filaments oscillate more. On the other hand, ejections are less obstructed by the filaments and thus dominate the inner-outer flow interactions.
@@
In order to isolate the effects of the filament flapping on turbulence, we have “frozen” them in one of their instantaneously deflected configurations and allowed the flow to develop. The turbulent shear is depleted with respect to the case in which the filaments are free to flap, since less energy is “pumped” into the turbulent fluctuations by their motion. Furthermore, the “frozen” canopies also prove more effective in shielding the inner flow. Consequently, throughout all cases, the drag is reduced with respect to the flexible ones upon “freezing” the canopy, and the maximum reduction is achieved for intermediate values of the flexibility, where the filaments exhibit large amplitude, regular fluctuations. Consistently with our previous work on the flapping states of a clamped flexible fiber in wall turbulence (Foggi Rota et al., 2023), also here we observe two regimes of motion for the filaments: one dominated by their structural natural response and one by turbulence. At the transition between the two, resonance occurs. Furthermore, in the turbulence dominated regime, all the filaments exhibit a dominant flapping frequency of abut $0.5U_{b}/H$ in the spanwise direction regardless of their structural parameters, once again supporting the outcome of our precursory study. Nevertheless, here, an increase in the density of the filaments drives them to exhibit their natural response since they better shield each other and get less deflected by the mean flow. Separately observing the terms of structural energy balance we also notice that the elastic energy peaks well within the regime dominated by the natural response, differently from the case of free fibers in homogeneous isotropic turbulence (Rosti et al., 2018a), where the peak is found at the transition between the two regimes. Furthermore, here, the kinetic energy saturates as soon as the filaments approach the turbulence dominated regime.
@@
[Acknowledgements] The research was supported by the Okinawa Institute of Science and Technology Graduate University (OIST) with subsidy funding from the Cabinet Office, Government of Japan. The authors acknowledge the computer time provided by the Scientific Computing section of Research Support Division at OIST, and by the HPCI System Research Project with grants hp210025 and hp220402.

### 14 May 2024 v2

@@ 3.6 Filament density variation
In order to investigate the effects of a variation in the inertial properties of the filaments, we consider two additional values of the density ratio $\{1.0+1.46\cdot 10^{-2},1.0+1.46\cdot 10^{-1}\}$ starting from the case at $Ca=25$ and $\rho_{s}/\rho_{f}=1.0+1.46\cdot 10^{-3}$.
@@ 3.7 The effect of the filaments’ flexibility
The drag reduction observed freezing a flexible canopy in its instantaneous configuration can therefore be imputed to an overall reduction of the turbulent activity with respect to the flexible case. We identify two phenomena contributing to such effect: first, upon freezing the filaments, no more energy is pumped into the small scale turbulent fluctuations of the flow by their flapping motion. This is consistent with the maximum drag reduction being attained in the intermediate regime ($Ca=25\sim 50$), where the large amplitude flapping of the filaments is most regular. As further discussed in the following section, in facts, at intermediate values of $Ca$ the natural response of the filaments resonates with a characteristic frequency of the turbulent forcing. Second, the frozen canopies are more effective then the flexible ones at preventing intense vertical interactions between the inner and the outer flows, thus yielding a weakly driven turbulent state within the canopy and depleting the turbulent fluctuations close to the shear layer.
@@ 4.1 Characterisation of motion
Multiple quantities can be observed in order to characterise the motion of the flexible filaments constituting the canopy. Here, we focus on the spanwise velocity of the Lagrangian points at the tip of the filaments: we prefer the measurement of a velocity to that of a displacement since it is more directly related to the turbulent fluctuations of the flow, and we favour the spanwise direction as it is the one less affected by the constraints of the problem. The streamwise and vertical dynamics of the filaments are the most influenced by the constraints, i.e., the inextensibility and the presence of the wall. We therefore collect the Lagrangian velocity signal in the spanwise direction at the tip of all the filaments throughout all the values of $Ca$ considered in this study, and highlight its features both in the time and frequency domains. For the purpose of visualisation, we also choose one canopy filament per value of $Ca$ and show in panel $a$ of figure 18 the velocity signal over ten bulk time units extracted after the attainment of a fully developed flow state. Here we also discuss the results of the simulation at $Ca=500$ to encompass the case in which the filaments are fully compliant to the flow.
@@
We immediately observe that the amplitude of the fluctuations increases with the flexibility of the filaments up to $Ca=25\sim 50$, where it appears to saturate. Furthermore, the small amplitude fluctuations of the most rigid case ($Ca=1$) are characterised by a nearly sinusoidal shape, modulated in amplitude by slower dynamics. Such behaviour contrasts with the motion of the most flexible filaments ($Ca=100\sim 500$), which instead exhibit fluctuations with a less definite shape and period. In these cases it is also possible to appreciate abrupt and fast small scale fluctuations on top of the slower dynamics: those arise from the collisions of the filament with the wall, and indeed are not observed for the lowest values of $Ca$ as the filament never gets sufficiently deflected. In between these two behaviours ($Ca=50$), the large amplitude fluctuations are almost sinusoidal in shape and exhibit a more definite period; they therefore appear compatible with a resonance between the natural structural response of the filaments (responsible for the nearly sinusoidal oscillations of the most rigid cases) and the turbulent fluctuations of the flow (dominating the motion of the filaments in the most flexible cases).
@@
To confirm the picture above, we investigate the energy spectrum of the signal, $E_{z}^{tip}=\frac{1}{2}\overline{\overline{\mathcal{F}(w_{tip})\mathcal{F}^{*}(w_{tip})}}$, where $\mathcal{F}$ is the temporal Fourier’s transform operator, * is the conjugate of an arbitrary complex number and the double over bar denotes ensemble-averaging on all the filaments of each case. After computing $E_{z}^{tip}$ over a sufficiently long time span cleaned of any initial transient, we report it in panel $b$ of figure 18 for the different values of $Ca$ considered. We also mark with vertical dashed lines the natural frequencies associated to the first structural mode of the different filaments, $f_{nat}$, computed as described in §2. For the lowest values of $Ca$ ($1$ to $50$), the spectrum peaks in correspondence of the natural frequency (and higher harmonics), thus confirming that the structural natural response dominates the dynamics of the most rigid filaments. The value of $f_{nat}$ decreases with $\gamma$, hence it gets shifted towards the left for increasing values of $Ca$. The peak of the spectrum also obeys this shift in frequency up to $Ca=50$, while beyond that its position does not to change significantly with the $Ca$. Such behaviour suggests that the most flexible filaments adjust to a motion no more dictated by their structural properties, but presumably compliant to the turbulent fluctuations of the flow. The transition between the two distinct regimes of motion occurs at $Ca\approx 50$, and thus we justify the large amplitude nearly sinusoidal oscillations observed in that case with the resonance between the first structural mode of the filaments and a characteristic frequency of the turbulent forcing to which they are exposed.
@@
The identification of two distinct regimes of motion is not unexpected: previous investigations addressing the motion of flexible fibres in homogeneous isotropic turbulence (Rosti et al., 2018a, 2020; Olivieri et al., 2022b) highlighted the existence of a structure-dominated and a turbulence-dominated regime, partially similar to what found when studying vortex induced vibrations (Bearman, 1984). We recently characterised numerically the motion of a flexible fibre clamped in wall turbulence over a wide range of structural parameters (Foggi Rota et al., 2024), also observing the emergence of the two regimes discussed above. Recently, Fu et al. (2023) experimentally investigated the individual motion of an isolated flexible plant model and how that changes when multiple plants are arranged in a canopy, while Monti et al. (2023) studied the motion of the filaments in a flexible canopy, mostly to characterise their collective dynamics (honami/monami). To reconcile the different insights offered by all these works, in the following section we further describe the flapping state of the individual filaments and better contextualise our observations.
@@ 4.2 The flapping state
To better investigate the different dynamical regimes of the filaments, we compute their dominant flapping frequency along the spanwise direction, $f^{flap}_{z}$, i.e., the peak location of the spectra reported in panel $b$ of figure 18, and plot it against the natural frequency associated to the corresponding value of $Ca$, i.e., with reference to section §2, $f_{nat}\approx\frac{3.516}{dh^{2}}\sqrt{\frac{\rho_{f}dh^{3}U_{b}^{2}}{2Ca\rho_{s}\pi^{3}}}\approx 0.45\frac{U_{b}}{h}\sqrt{\frac{\rho_{f}}{\rho_{s}}\frac{h}{d}\frac{1}{Ca}}$ Since for the highest values of $Ca$ the position of the spectral peak does not appear to undergo any significant change with the structural parameters of the filaments, we denote it as $f_{turb}$ (i.e., the characteristic flapping frequency of the filaments in the turbulence dominated regime) and employ it to adimensionalise both $f^{flap}_{z}$ and $f_{turb}$. The outcome is reported in panel $a$ of figure 19, where flapping frequencies close to the natural one lay on the dashed-dotted line with unitary slope, $f^{flap}_{z}/f_{turb}=f_{nat}/f_{turb}$, while those close to the chosen value of $f_{turb}$ approach the dashed horizontal asymptote. We thus confirm that the most rigid filaments exhibit their natural response, while the most flexible ones deviate from it and approach $f^{flap}_{z}\approx f_{turb}=0.5U_{b}/H$. Such frequency, independent from the structural characteristics of the filaments, is consistent with the outcome of our previous investigation (Foggi Rota et al., 2024) focused on the flapping states of an isolated flexible fibre clamped in wall turbulence. Also in that case indeed, we identified a turbulence-dominated regime for the highest values of $Ca$ where the fibre is compliant to the flow and its motion is characterised by a slow flapping with a period comparable to the bulk time-scale of the flow. Furthermore, despite the absence of a physical justification for $f^{flap}_{z}\approx 0.5U_{b}/H$, we observe that its occurrence in both the motion of an isolated flexible fibre in a full-channel flow and in the motion of the canopy filaments in the the half-channel considered here supports the reproducibility and generality of the result.
@@
Figure 19: Flapping states of the filaments for different values of $Ca$, hence $f_{nat}/f_{turb}$. We first show (panel $a$) the dominant frequency of oscillation along the spanwise direction, $f^{flap}_{z}$, extracted as the peak location of the spectra in figure 18. Squares denote data from the cases at different density ratios. (panel $b$) We also report the trends of the elastic ($\Gamma_{s}$) and kinetic ($K_{s}$) energy of the filaments per unit length, averaged over time and over all the filaments. The error-bars are computed carrying out the same measurement on the time signals truncated to their first half, and fall within the markers size for most points.
@@
In order to assess the effect of an increase in the structural density of the filaments on their dynamics, we also report in the map of figure 19, panel $a$, the results of the cases at ${\rho_{s}}/{\rho_{f}}=\{1+1.46\cdot 10^{-2},1+1.46\cdot 10^{-1}\}$, respectively associated to a higher and a lower value of the natural frequency. Differently from what found in our previous investigation of a single filament (Foggi Rota et al., 2024), here an increase in ${\rho_{s}}/{\rho_{f}}$ prevents the filaments from flapping at $f_{turb}$ and they thus exhibit their natural response. This difference originates from the fact that denser filaments yield a stronger obstruction of the mean flow in the case of a canopy (as elucidated in §3.6), and consequently get less deflected and maintain an almost upright configuration, swaying at $f_{nat}$. This effect is lost in the case of an isolated fibre, as the back-reaction on the mean flow is negligible; in this isolated case the fibre is less shielded and consequently gets more deflected by the flow, swaying with it.
@@
The effect of the clamp on the dynamics of the filaments is further understood observing the variation of their elastic energy with $Ca$, while their motion is better described in terms of kinetic energy. To those ends we therefore consider the weak form of equations 3 and 4, integrated over a time-span $\tilde{t}$ and over the length of the filaments,
@@
represents the structural kinetic energy per unit length, likewise averaged. We report the trends of $\Gamma_{s}$ and $K_{s}$ with $f_{nat}/f_{turb}$, and hence with the $Ca$, in panel $b$ of figure 19. While the magnitude of $\Gamma_{s}$ is comparable to that measured by Rosti et al. (2018a), the position of its peak contrasts with their findings. Here, in facts, $\Gamma_{s}$ peaks well within the regime dominated by the natural response of the filaments, confirming what can also be observed in the case of an isolated clamped fibre. We therefore deduce that the clamp to the wall, constraining the shape of the filaments, hinders their deflection and shifts the maximum of $\Gamma_{s}$ towards higher values of $\gamma$. The kinetic energy, instead, follows a trend consistent with the amplitude of the fluctuations in figure 18, panel $a$, increasing with the $Ca$ in the most rigid cases and reaching a plateau close to the resonance between $f_{nat}$ and $f_{turb}$.
Characterising the dominant dynamics of our canopy filaments, we have been able to compare our results with those attained considering a single fibre clamped in wall turbulence (Foggi Rota et al., 2024) and plants with a different geometry (Fu et al., 2023). We have also isolated the role played by the sheltering effect of the canopy upon increasing the density of the filaments, and by the model adopted to describe them. Nevertheless, we emphasise that the present investigation is unable to capture any sub-dominant dynamics of the filaments, such as the contribution of turbulence to their motion in the regime dominated by the natural response. Regardless of the value of $Ca$, in facts, the coherent motion of the filaments (honami/monami) investigated by Monti et al. (2023) remains dictated by the turbulent structures in the flow.
@@ 5 Conclusions
Filaments of increasing flexibility are more deflected by the flow, thus reducing the frontal area of the canopy, and they yield a weaker drag discontinuity at the canopy tip, associated to velocity fluctuations less intense then in the rigid case. Consequently, the canopy drag decreases increasing the flexibility up to saturation, attained when the fully deflected filaments start piling on top of each other parallel to the wall. Our data compare well to the experimental measurements of Shimizu et al. (1992) and Ghisalberti & Nepf (2006). The spectra of the turbulent kinetic energy within the canopy confirm the spectral short cut mechanism described by Finnigan (2000) and Olivieri et al. (2020), while a more regular energy cascade is attained above the canopy tip. Furthermore, the intensity of the velocity fluctuations within the canopy increases with its flexibility due to the increased motion of the filaments. Outside the canopy, instead, the opposite trend is attained by virtue of the generation of a stronger shear layer in the more rigid cases.
The characterisation of the turbulent state at different distances from the bottom wall suggests a multi-layer approach to turbulence modelling in canopy flows. Within the canopy, turbulence is quasi 2D close to the bottom wall and it approaches an isotropic state between the inner inflection point and the virtual origin of the outer flow. The picture becomes more blurred increasing the flexibility, as the flow in the canopy turns less isotropic. A strongly anisotropic state is attained approaching the canopy tip, and the flow exhibits a behaviour consistent with what observed in conventional channel flows moving above it, thus supporting outer similarity arguments. Turbulence inside the canopy appears to be sustained by intense events generated in the outer flow, induced by the structures populating the close proximity of the shear layer. Those structures cause sweeps and ejections dominating the interaction between the inner and the outer flows, but only the strongest sweep events are able to penetrate the canopy due to the obstruction exerted by the deflected filaments. The effect increases with the flexibility, yielding a more weakly driven turbulent state within the most compliant canopies even though the filaments oscillate more. On the other hand, ejections are less obstructed by the filaments and thus dominate the inner-outer flow interactions.
@@
In order to isolate the effects of the filament flapping on turbulence, we have “frozen” them in one of their instantaneously deflected configurations and allowed the flow to develop. The turbulent shear is depleted with respect to the case in which the filaments are free to flap, since less energy is “pumped” into the turbulent fluctuations by their motion. Furthermore, the “frozen” canopies also prove more effective in shielding the inner flow. Consequently, throughout all cases, the drag is reduced with respect to the flexible ones upon “freezing” the canopy, and the maximum reduction is achieved for intermediate values of the flexibility, where the filaments exhibit large amplitude, regular fluctuations. Consistently with our previous work on the flapping states of a clamped flexible fiber in wall turbulence (Foggi Rota et al., 2024), also here we observe two regimes of motion for the filaments: one dominated by their structural natural response and one by turbulence. At the transition between the two, resonance occurs. Furthermore, in the turbulence dominated regime, all the filaments exhibit a dominant flapping frequency of abut $0.5U_{b}/H$ in the spanwise direction regardless of their structural parameters, once again supporting the outcome of our precursory study. Nevertheless, here, an increase in the density of the filaments drives them to exhibit their natural response since they better shield each other and get less deflected by the mean flow. Separately observing the terms of the structural energy balance we also notice that the elastic energy peaks well within the regime dominated by the natural response, differently from the case of free fibers in homogeneous isotropic turbulence (Rosti et al., 2018a), where the peak is found at the transition between the two regimes. Furthermore, here, the kinetic energy saturates as soon as the filaments approach the turbulence dominated regime.
@@
[Acknowledgements] The research was supported by the Okinawa Institute of Science and Technology Graduate University (OIST) with subsidy funding from the Cabinet Office, Government of Japan. The authors acknowledge the computer time provided by the Scientific Computing section of the Core Facilities at OIST, and by the HPCI System Research Project with grants hp210025 and hp220402.
@@
[Data availability statement]The plotted data are available at https://groups.oist.jp/cffu/foggirota2024jfluidmech.
[…]
[Author ORCIDs]
Giulio Foggi Rota ID : https://orcid.org/$0$-0002-4361-6521
Alessandro Monti ID : https://orcid.org/$0$-0003-2231-2796
Stefano Olivieri ID : https://orcid.org/$0$-0002-7795-6620
Marco Edoardo Rosti ID : https://orcid.org/$0$-0002-9004-2292
Appendix A Flow visualisations
We devote this appendix to the visualisation of the instantaneous flows fields. Panel $a$ of each figure reports a slice in the $x-y$ plane sampled from the rigid canopy case, while panel $b$ shows a similar slice from the flexible canopy case, at $Ca=100$. The flow is oriented from left to right. We display the instantaneous values of $u^{\prime}$ (figure A), $v^{\prime}$ (figure 20) and $p^{\prime}$ (figure 21): they indicate a less strongly driven turbulent state in the flexible canopy case, with enhanced spacial coherence compared to the rigid one. Furthermore, areas of markedly low pressure can be seen at the canopy tip, denoting the cores of the spanwise rollers populating that region. In figure 22 we also visualise the instantaneous flow anisotropy. On average, the flow above the rigid canopy appears more anisotropic than that above the flexible one, while the opposite is true within the canopy. Such scenario is consistent with the plots reported in figure 10.
{bottomfigure}[Uncaptioned image]
a)[Uncaptioned image]b)We report vertical slices of the fluctuating streamwise velocity in and above a rigid (panel $a$) and a flexible (panel $b$) canopy with $Ca=100$, with the mean flow directed from left to right. A colour scale going from violet to orange is adopted, ranging in $[-0.8,0.8]U_{b}$.
a)Refer to captionb)
Figure 20: We report vertical slices of the fluctuating wall-normal velocity in and above a rigid (panel $a$) and a flexible (panel $b$) canopy with $Ca=100$, with the mean flow directed from left to right. A colour scale going from violet to orange is adopted, ranging in $[-0.6,0.6]U_{b}$.
a)Refer to captionb)
Figure 21: We report vertical slices of the fluctuating pressure in and above a rigid (panel $a$) and a flexible (panel $b$) canopy with $Ca=100$, with the mean flow directed from left to right. A colour scale going from violet to orange is adopted, ranging in $[-0.2,0.2]\rho U_{b}^{2}$.
a)Refer to captionb)
Figure 22: We report vertical slices of the instantaneous flow anisotropy in and above a rigid (panel $a$) and a flexible (panel $b$) canopy with $Ca=100$, with the mean flow directed from left to right. A colour scale going from violet to orange is adopted, ranging in $[0.0,0.58]$.